OpenStudy (darkigloo):

2H2 + O2 -> 2H2O How many grams of O2 and H2 were reacted to produce 30.0 g of H2O?

OpenStudy (darkigloo):

@hartnn

OpenStudy (darkigloo):

@iambatman @nincompoop @Loser66

OpenStudy (darkigloo):

@Preetha @hba @tHe_FiZiCx99 @aaronq @BioEpic @prowrestler

OpenStudy (anonymous):

\[2H_2+O_2\rightarrow 2H_2O\] As you can see, \(n_{H_2}=2n_{O_2}=n_{H_2O}\) First, calculate the amount of \(H_2O\). Now you get \(n_{H_2}=n_{H_2O}\), \(n_{O_2}=\frac{1}{2}n_{H_2O}\). You can easily calculate the mass of \(H_2\) and \(O_2\).

OpenStudy (darkigloo):

what do you mean by calculate the amount of H2O? Would that be 30g?

OpenStudy (anonymous):

It is the amount of \(H_2O\) in moles. \(n=\frac{m}{M}\)

OpenStudy (darkigloo):

so 1.7 moles of H2O

OpenStudy (anonymous):

Yes, about \(\frac{5}{3}\).

OpenStudy (darkigloo):

and now what do i do?

OpenStudy (anonymous):

\(n_{H_2}=n_{H_2O}\)

OpenStudy (darkigloo):

would it be 1.7 mol H2 and 0.85 mol O2? :/

OpenStudy (anonymous):

Correct. :)

OpenStudy (darkigloo):

so i would change that to grams by multiplying by the mass of H2 and O2?

OpenStudy (anonymous):

Yes, the molar mass M.

OpenStudy (darkigloo):

Would it be 1.7mol H2 x 1g = 1.7g H2 0.85mol O2 x 16g = 13.6g O2 or 1.7mol H2 x 2g = 3.4g H2 0.85mol O2 x 32g = 27.2g O2

OpenStudy (anonymous):

Actually the second one is correct. A hydrogen molecule is formed by two H atoms.

OpenStudy (darkigloo):

ok thank you.

OpenStudy (anonymous):

It's my pleasure :)

OpenStudy (preetha):

Good explanation there tinyworm!

OpenStudy (anonymous):

Thank you, @Preetha.